OpenStudy (anonymous):

solve for x...x/2-4=-5.....plz show the steps:)

OpenStudy (anonymous):

1.X/2-4=-5(add 4 to both sides), 2.X/2=-1(multiply the x/2 with the reciprocal so multiply both sides by 2), 3.X=-2